A college professor hands out a list of 1010 ​questions, 55 of which will appear on the final examination for the course. One of the students taking the course is pressed for time and can prepare for only 99 of the 1010 questions on the list. Suppose the professor chooses the 55 questions at random from the 1010. a. What is the probability that the student will be prepared for all 55 questions that appear on the final​ examination? what about 4 or 3 questions?
Answer #1

a)probability that the student will be prepared for all 5 questions that appear on the final​ examination

=P(selecting 5 from 9 he has prepared and none from he has not prepared)

=9C5*1C0/10C5 =126/252 =1/2 =0.5

b)P(student prepare for 4) ==9C4*1C1/10C5 =126*1/252 =0.5

P(student prepare for 3)=0   (as it is not possible cause in this case he should not be prepared for 2 questions while he is not prepared for 1 questions only)